OpenStudy (anonymous):

the average of three consecutive odd numbers is -3

OpenStudy (anonymous):

The average of 3 numbers would be the sum of them divided by 3 so (a+b+c)/3=-3 (eq. 1) and knowing they are consecutive odd numbers a=a (we don't know yet) b=a+2 (+2 because b and c have to be odd) c=b+2 so c=a+2+2 substituting eq.1 a+b+c=-9 a+(a+2)+(a+4)=-9 3a+6=-9 3a=-9-6=-15 a=-15/3=-5 so b = -5+2 = -3 and c = -5+4 = -1 proof (-5-3-1)/3=-3